OpenStudy (anonymous):

if F(x) = log small 2 base x / x^2 , find f(1/2) ?? help please!

OpenStudy (anonymous):

Is this the question? \[\frac{\log_{x} {2}}{x^2}\] or this \[\frac{\log_{2} {x}}{x^2}\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

help

OpenStudy (anonymous):

please

OpenStudy (lgbasallote):

\[\frac{\log_2 (\frac{1}{2})}{(1/2)^2}\]

OpenStudy (lgbasallote):

i just substituted 1/2 into x

OpenStudy (lgbasallote):

\[\frac{\log_2 2^{-1}}{\frac 14}\] can you continue?

OpenStudy (anonymous):

how did you get 2^-1?

OpenStudy (anonymous):

oh wait ok..

OpenStudy (anonymous):

can you continue

OpenStudy (anonymous):

? please